Need help with finance. Please please help. Thank you so much.
To what amount will the following investments accumulate?
a. $5,000 invested for 10 years at 10 percent compounded annually
b. $8,000 invested for 7 years at 8 percent compounded annually
C. $775 invested for 12 years at 12 percent compounded annually
D. $21,000 invested for 5 years at 5 percent compounded annually

Answer by rfer(12660)

FV=P(1+r)^n
--------------------------
a) 5000(1+.10)^10
5000(1.10)^10
5000*2.59734246=$12986.71
-------------------------
b) 8000(1+.08)^7
8000(1.08)^7
8000*1.713824269=$13710.59
---------------------------
The procedure is the same for the other two.
Give them a try.
Good luck.